# On the approximability of random-hypergraph MAX-3-XORSAT problems with quantum algorithms

### v1

@@ VI.2 Trial minimum annealing performance
We also tested the TMA formulation of spectral folding–the formulation for which we can make the most reliable analytical predictions–as plotted in Fig. 8. For these variations we used runtimes $t_{f}=N/12$ and $dt=0.025$; note that this choice of $t_{f}$ is a factor of $8/3$ larger than in our QAOA simulations but with the same scaling. In smaller system studies similar qualitative performance was observed for shorter $t_{f}$ (such as $N/24$ or $N/32$). For this variation we used a 3-XORSAT $H_{L}$–the formulation for which we could predict performance in Sec. IV–with the minimum energy set to $-2N$ via $C\left(t\right)$, which was linearly ramped down to zero by $t_{f}$, and simple sinusoidal ramp profiles with $t_{r}=N/24$; the transverse field strength $\kappa$ during the main evolution was 1.3. The careful reader may note that choosing $C\left(t\right)$ to set the minimum energy to $-2N$ instead of $-N$ is naively suboptimal, as all other things being equal sweeping over a larger energy range increases $W$ in Eq. (29), and should reduce the returned probabilities $P_{q}\left(N\right)$ by an appropriate prefactor. However in our simulations this choice consistently improved both the prefactors and scaling, e.g. the value of $q_{a}$, as compared to choosing a minimum energy $-N$ for $H_{L}$. We suspect this has to do with the band structure considerations described in Appendix B but due to the complexity of the problem, are unable to make a quantitative prediction.
@@
The performance of the two approaches is qualitatively similar with subtle differences as we vary the returned approximation ratio $q$. At lower approximation ratios the AQC formulation returns higher probabilities, at lower total gate count since there are no ramping steps and no additional gates associated with adding $H_{L}$. However, at higher approximation ratios the TMA formulation appears to be better able to approach the approximation target $A=0.85$. As discussed in the algorithm definition, folded optimization will definitionally fail to consistently return energies significantly below $AE_{\mathrm{GS}}$, and we expect it to break down as $A$ gets too close to 1 given that QAOA and similar methods fail to reliably approximate these problems. Choosing the best value for $A$ is thus a subtle issue that depends on the problem class; for extensions of this method to hard CSPs it will necessarily change from one problem class to the next.
@@
Likewise, as mentioned in section IV.5, one can replace the random 3-XORSAT lowering Hamiltonian in TMA for a simpler set of linear $Z$ biases, reducing the gate count per timestep and, potentially, increasing the per-state tunneling rate by implementing a shallower cost-per-flip curve. In comparison to the 3-XORSAT variation discussed in the previous paragraph, to achieve good performance we needed to double the ramp time. This protocol seemed to be more sensitive to performance degradation from heating during ramps. The algorithm also benefitted from adjusting $C\left(t\right)$ so that the minimum energy of $H_{L}$ was $-3N$, as compared to $-2N$ for the 3-XORSAT $H_{L}$. Relative performance for $A=0.75$ is comparable to the other variations, as illustrated in Fig. 9, though the individual $P_{q}\left(N\right)$ show more significant non-monotonicity that makes reliable curve fitting challenging. This issue is even more pronounced for $A=0.85$ (data not shown), to the point that we did not quote $q_{a}$ values for that variation in the Table 1.
@@
VI.3 Discussion of our spectrally folded optimization results
@@
The reader may note that all of these simulations targeted energies in the range $A=0.75$ to $A=0.85$, well above the theoretical prediction of $\sim 0.6$. We also tested $A=0.65$ at smaller scales (data not shown) for both formulations but found no region of the parameter space where it showed meaningful improvements over choosing $A=0.75$ or higher. When increasing $A$ to 0.85 we were not able to meet this target with constant probability, though for some cases we did see improvements in the returned achievable approximation ratio $q_{a}$ when compared to $A=0.75$. We did not run tests with $A>0.85$ due to system size constraints, though comparing performance between $A=0.75$ and $A=0.85$ in table 1 suggests we have found the performance ceiling for these methods. Given the comparatively small system sizes available to classical simulation our ability to draw meaningful scaling distinctions with very small changes in $A$ is limited. Considering the results of all our numerical simulations, the minimum achievable approximation ratio for our PPSP problem classes is at least 0.7 (often 0.75) for the quadratic AQC variation, and at least 0.75 (often 0.8) for the 3-XORSAT TMA variation.
@@
As summarized in table 1, the contrast between the clear super-polynomial decay of higher approximation ratios with classical methods and QAOA (Figs. 4-6), and the constant probabilities returned by folded quantum optimization (Figs. 7-9) is stark. In particular, for the established methods, the polynomial time approximation threshold is set by the problem structure, specifically the relative energy of the high energy clustering phase, and as our PPSPs become more extremal this threshold steadily worsens as a fraction of the energy of the planted ground state. We found no evidence for an exponential separation in approximation power between local classical searches and QAOA/AQC. At lower constraint densities, the clustering energy is not far above the ground state and spectrally folded quantum optimization provides no benefits for approximation, though it can still show scaling advantages for returning states close in Hamming distance to the ground state (see appendices). At higher constraint densities however, spectral folding is able to return states close to $G$, in both energy and Hamming distance, consistently and without degradation as $N_{C}/N$ increases. The performance of all spectral folding variations tested was broadly similar, with the 3-XORSAT TMA variation returning the highest approximation ratio but at the highest prefactor cost in gate count.
@@
This illustrates the fundamentally different structure of collective tunneling in this approach. Where high depth QAOA, AQC and other direct methods attempt to find the ground state, asymptotically fail (given exponential scaling), and return approximate states passed after this missed opportunity, spectrally folded optimization deforms the cost function to search for states in an exponentially large hyperspherical shell, avoiding the interference and weakening field issues that are qualitatively responsible for QAOA’s lack of obvious quantum advantage. In the interest of fair comparisons the total runtime per-shot of all routines is $O\left(N\right)$ Hamiltonian calls. With this simple linear scaling we are able to ensure that any exponential scaling of the time to solution not visible from our figures or fits must involve very small exponents, though of course we cannot rule out such behavior on numerics alone. We thus do not see our numerical simulations as a precision scaling benchmark of folded quantum optimization, but rather a test of the basic veracity of our conservative analytical predictions in section IV. The results of our simulations, for a wide range of parameters, all suggest that our theory is sound.
VI.4 Extensions of these results
@@
Having thoroughly explored MAX-3-XORSAT in this work, it would be interesting to test spectral folding methods on other hard CSPs or problems that can be straightforwardly formulated as such. We expect that the derivation of the achievable approximation ratio through the overlap of dressed states could generalize with some modifications to many other problem classes. We focused on MAX-3-XORSAT due to the simplicity of its structure, classical approximation hardness, and the fact that its exponential difficulty scaling is obvious at small $N$ for more standard classical and quantum approaches.
@@
One interesting potential exception to this argument is quantum Monte Carlo. Being a stoquastic problem, a folded quantum spin glass can in principle be efficiently simulated using QMC Isakov et al. (2016); Andriyash and Amin (2017); Jiang et al. (2017a, b); King et al. (2019), which for a uniform field has in some cases proven to be an effective quantum-inspired classical solver Heim et al. (2015); Albash and Lidar (2018). Incorporating spectral folding into these algorithms is possible, though the loss of locality makes evaluating each update much more expensive. As QMC is, fundamentally, an energy-based classical local update rule combined with many replicas, we do not expect it to overcome the entropic barrier in MAX-3-XORSAT the way that true quantum evolution can. [^7] As remarked earlier, we found the scaling of linear spectral folding in the AQC framework, which has no influence on the high energy spectrum at all, to be close to that of quadratic folding for the same schedule. As finding the deep minimum is an entropic barrier challenge for classical algorithms, we do not expect simulated quantum annealing through QMC to be able to replicate the true quantum result. All that said, there may be some narrow cases where incorporating spectral folding in a QMC calculation could prove useful as a classical solver; this would be an interesting avenue for future research. If it turns out that many or all of these instances are efficiently approximable classically through QMC with a folded spectrum, that would be a very significant discovery in its own right.
@@ VII Conclusions and Outlook
Using the NP-hard MAX-3-XORSAT problem class, we explored the question of classical and quantum approximation hardness from a practical, mechanism-focused point of view. Guided by theoretical intuition, we proposed a class of instances called planted partial solution problems (PPSPs) which we showed are empirically hard for both exact and approximate optimization for classical searches and established quantum methods such as the quantum approximate optimization algorithm (QAOA) and adiabatic quantum computing (AQC). Through extrapolation and qualitative analysis of a rigorous calculation of the typical-case minimum gap at the first order paramagnet to spin glass transition bottlenecking MAX-3-XORSAT, we were able to identify two effects that significantly impede high-depth QAOA’s ability to approximate the hardest instances, namely weakening transverse field and destructive interference. We then proposed a novel algorithmic update, called spectral folding, that does not suffer from these two issues. Spectral folding is conceptually simple; the analytical analysis of its performance is perhaps less so. And rather intriguingly, it works by applying classical modifications to the classical cost function being optimized, which only provide benefits for quantum optimization.
@@
Using a resummed extensive-order perturbation theory, we were able to predict a constant fraction approximation guarantee for our difficult random hypergraph PPSPs, and consequently, an exponential quantum speedup in the classically hard regime. To further support our claims, we performed a series of numerical simulations of high-depth QAOA, classical optimization routines, and spectral folding methods, out to the largest sizes that we could feasibly reach while still being able to gather good statistics. These numerical results support our claims and we did not discover any meaningful discrepancies or red flags in them; every major prediction here has been numerically checked to the extent we found reasonably possible with current supercomputing resources.
@@
The implications of an effective fast approximation guarantee through spectral folding–or any quantum algorithm, for that matter–are profound, and even restricting ourselves to MAX-3-XORSAT it is important to ask what types of hypergraph might cause it to fail. To be clear, “failure” in this case means that, for a class of hypergraphs, some property invalidates the analytical predictions we made and restricts polynomial-time approximation to values of $A$ near zero. If a subsequent, more sophisticated analysis shows, for example, that due to some subtle effect missed in our resummed extensive order perturbation theory, the variations we propose are asymptotically limited to $A=0.4$ for random hypergraphs instead of our prediction of $\geq 0.6$, we would not consider that a general failure of the algorithm. Given the exponential density of target states, for folded quantum optimization to fail the per-state tunneling rate needs to be much worse than what our calculations and simulations return; even doubling the per-state tunneling decay exponent in Eq. (IV.4) yields $A\simeq 0.18$, a much worse approximation ratio than we predict and observe in simulations, but still a constant fraction better than random guessing.
@@
We are not so hubristic as to claim it is impossible that there is some hidden effect that substantially worsens scaling at large $N$ for random hypergraphs, which is not captured by our theory and is invisible in our numerics. But we see no evidence for it, have no idea what such an issue could be, and formulated our theory such that the simplifying approximations we made were more likely to underestimate the achievable approximation ratio $A$ than overestimate it. Excellent empirical performance in simulation supports this interpretation. We see it as much more likely that one can structure PPSP hypergraphs in some non-random way as to violate the core assumptions of our calculations and become inapproximable. The conditions on such hypergraphs are, however, fairly strict; besides needing to more than double the exponential decay rate of per-state tunneling as compared to random hypergraphs, whatever property is responsible for the slowdown must be resilient both to the addition of a random, uncorrelated problem with similar constraint density and ground state energy, and to spectral deformations such as quadratic folding. Further, such graphs must be relatively dense, as sparse problems are easy to approximate by solving random sub-problems, and their construction rules must not inadvertently render them amenable to classical optimization.
@@
Identifying hypergraph properties that cause this entire method to break down could lead to valuable new discoveries about macroscopic quantum tunneling physics and problem hardness, and further algorithm innovations in the steps needed to mitigate their effects. As we stated in the introduction, the underlying reasons for classical approximation hardness are generic and intuitive if often only applicable to extremal problem instances, but their quantum equivalents are not, and much more opaque. The correctness of our predictions here would imply that quantum approximation hardness may not be generic at all, and instead specific to as-yet undiscovered sets of problem properties. Let’s go exploring.

### v3

@@ VI.2 Trial minimum annealing performance
We also tested the TMA formulation of spectral folding, with results plotted in Fig. 11. For these variations we used runtimes $t_{F}=N/12$ and $dt=0.025$; note that this choice of $t_{F}$ is a factor of $8/3$ larger than in our TAQC simulations but with the same scaling. In smaller system studies similar qualitative performance was observed for shorter $t_{F}$ (such as $N/24$ or $N/32$). Following the principles in section II.5, these longer runtimes were chosen to ensure we did not need to worry about performance degradation due to heating from the ramps themselves. For this variation we used a 3-XORSAT $H_{L}$–the formulation for which we could predict performance in Sec. IV–with the minimum energy set to $-2N$ via $C\left(t\right)$, which was linearly ramped down to zero by $t_{f}$, and simple sinusoidal ramp profiles with $t_{r}=N/24$; the transverse field strength $\kappa$ during the main evolution was 1.3.
The careful reader may note that choosing $C\left(t\right)$ to set the minimum energy to $-2N$ instead of $-N$ is naively suboptimal, as all other things being equal sweeping over a larger energy range increases $W$ in Eq. (2), and should reduce the returned probabilities $P_{q}\left(N\right)$ by an appropriate prefactor. However in our simulations this choice consistently improved both the prefactors and scaling, e.g. the value of $q_{a}$, as compared to choosing a minimum energy $-N$ for $H_{L}$. We suspect this has to do with the band structure considerations described in Appendix C but due to the complexity of the problem, are unable to make a quantitative prediction.
@@
The performance of the two approaches is qualitatively similar with subtle differences as we vary the returned approximation ratio $q$. At lower approximation ratios the AQC formulation returns higher probabilities, at lower total gate count since there are no ramping steps and no additional gates associated with adding $H_{L}$. However, at higher approximation ratios the TMA formulation appears to be better able to approach the approximation target $A=0.85$. As discussed in the algorithm definition, folded optimization will definitionally fail to consistently return energies significantly below $AE_{\mathrm{GS}}$, and absent warping we expect it to break down as $A$ gets too close to 1 given that TAQC and similar methods fail to reliably approximate these problems. Choosing the best value for $A$ is thus a subtle issue that depends on the problem class; for extensions of this method to hard CSPs it will necessarily change from one problem class to the next.
@@
Likewise, as mentioned in section IV.3.3, one can replace the random 3-XORSAT lowering Hamiltonian in TMA for a simpler set of linear $Z$ biases, reducing the gate count per timestep and, potentially, increasing the per-state tunneling rate by implementing a shallower cost-per-flip curve. In comparison to the 3-XORSAT variation discussed in the previous paragraph, to achieve good performance we needed to double the ramp time. This protocol seemed to be more sensitive to performance degradation from heating during ramps. The algorithm also benefitted from adjusting $C\left(t\right)$ so that the minimum energy of $H_{L}$ was $-3N$, as compared to $-2N$ for the 3-XORSAT $H_{L}$. Relative performance for $A=0.75$ is comparable to the other variations, as illustrated in Fig. 12, though the individual $P_{q}\left(N\right)$ show more significant non-monotonicity that makes reliable curve fitting challenging. This issue is even more pronounced for $A=0.85$ (data not shown), to the point that we did not quote $q_{a}$ values for that variation in the Table 1.
@@
VI.3 Multi-stage filtered optimization performance
Figure 13: Probability of finding the ground state with multi-stage filtered optimization (top, approximately constant data), as compared to Trotterized adiabatic quantum evolution (bottom, exponentially decaying data), for $N_{C}=\left\{2,4,6,3\sqrt{N}/2\right\}N$ (blue, gold, green and red points), demonstrating the rough independence of constraint density–and, critically, the scaling of $P_{GS}\left(N\right)$ for the runtime specified–predicted through many-body theory in the text. The TAQC data, which all shows clear exponential decay, is taken from FIG. 9 and rescaled by a multiplicative factor so that the total per-shot quantum evolution time is the same for all data. This runtime is sub-quadratic (scaling approximately as $N^{1.5}$) and given by Eq. 46. For $N_{C}=2N$ we observe performance degradation at larger $N$, which turns out to be due to Trotter error heating; reducing $dt$ from 0.025 to 0.02 restores constant $P_{GS}\left(N\right)$ (purple triangles). We attribute the larger sensitivity to Trotter error at low $d_{C}$ to the increased local variance of the density of constraints, and it may be the case that $dt$ needs exhibit slow polynomial decay with $N$ to maintain asymptotically constant scaling; both issues are discussed in section IV.2. Black dashed lines plot the average over constraint densities for the two methods. Each MSFO datapoint is the average of 480 randomly generated instances.
For our final set of simulations, we studied the multi-stage filtered optimization of section IV.2, which for $d_{C}\gg 1$ and $\epsilon\ll 1$ is predicted to find the ground state in sub-quadratic total evolution time. In that section, we made two claims, the first concerning the stability of the transition point and the second concerning the validity of mean-field methods for predicting the minimum gap. If the second claim in particular asymptotically holds then the time to solution must scale polynomially, roughly as $N^{1.46}$ for the parameters ($x=2$ and $w=0.6$ in Eq. 12) chosen. Our goal in these sections was to test the basic veracity of these claims. In particular, we wanted to verify that for a polynomial quantum evolution time $t_{F}\left(N\right)$, predicted from a mean field analysis, the probability of finding the ground state $P_{GS}\left(N\right)$ would be asymptotically constant with increasing $N$.
To determine $t_{F}\left(N\right)$, we needed to take into account a few effects that are not relevant in the large $N$ limit. First, as discussed in appendix IV.1, the average cost per flip away from $G$ has meaningful small-$N$ deviations from the asymptotic result in Eq. 18. To take this into account, we numerically calculated the average $E\left(x\right)$ for $x$ flips away from $G$ and used that result as the cost function in computing $\Omega_{\mathrm{mf}}\left(N\right)$. For $x=2$, $w=0.6$ and $N\leq 30$, we obtained $\Omega_{\mathrm{mf}}\left(N\right)\propto N^{-0.16}$. Second, the slope $W\left(N\right)$ was not perfectly linear, but instead best fit by $W\left(N\right)\propto\sqrt{N^{2}+w_{0}^{2}}$. Third, let the avoided crossing occur at $s_{c}\left(N\right)$ (for total Hamiltonian $H=\kappa H_{D}+sH_{\mathrm{cost}}^{x,w}$). It turns out for $N$ in the range accessible for classical simulation that the scaling of the gap with $s$ is substantially asymmetric around $s_{c}$ (with a much sharper slope with $s>s_{c}$), and $s_{c}\left(N\right)$ likewise empirically varies as $s_{c}\left(N\right)=s_{0}+a/N$ for a constant $a$ and $s_{0}\simeq 0.3$ is the asymptotic result. Putting all these effects together, we posited a runtime scaling
$t_{F}\left(N\right)\propto\frac{\sqrt{N^{2}+w_{0}^{2}}}{\Omega_{\mathrm{mf}}\left(N\right)^{2}s_{c}\left(N\right)}.$ (46)
This function approximately scales as $N^{1.5}$ for the parameters chosen here, and was used as the input to set the runtime of our simulations. We set $dt=0.025$ to minimize heating from Trotter error (though as seen in figure 13, we needed to reduce it further for $d_{C}=2$), and let the runtime of the first stage be $t_{F}\left(N\right)$ and the second and third stages be $t_{F}\left(N\right)/2$ for each. Finally, we chose the prefactor of $t_{F}\left(N\right)$ so that $t_{F}\left(8\right)=1/4$, matching the runtime used in the TAQC simulations of section V.3.
The results of our simulations are plotted in FIG. 13, and support our assertion of a polynomially scaling time to find the true ground state for PPSPs in the approximation-hard regime. With suitably small $dt$ we find that $P_{GS}\left(N\right)$ is asymptotically independent of constraint density and asymptotically constant with $N$, at least up to $N=25$, the largest sizes simulated. Both the constant scaling, and the approximate $d_{C}$ independence, are implied by our theory. The performance advantage of MSFO is particularly stark when compared to TAQC, as seen in the figure. Per Eq. 2 we assume that the TAQC success probability scales roughly as $P_{GS}\propto\Omega_{0}\left(N\right)^{2}t_{F}\left(N\right)/W\left(N\right)$, and therefore to make the comparison fair we took the probabilities computed in section V.3 and multiplied them by the appropriate factor $\left(8/N\right)\left(t_{F}\left(N\right)/t_{F}\left(8\right)\right)$ to afford them the same total quantum evolution time as MSFO.
We find the result for $d_{C}=2$ to be particularly surprising, in some sense, given that the local constraint density variation is proportionally large in that case, and that the clustering energy is 75-80% of $E_{GS}$ so we do not think transverse field chaos can be a priori ignored once filtering is applied. The only algorithm modification we needed to make to match the performance at higher constraint densities was to reduce $dt$ (and thus increase the circuit depth, so the total evolution time was the same) by $20\%$. We attribute the increased sensitivity to Trotter error to the larger local variations in the cost per flip of the classical $H_{P}$; recall that we have normalized each problem so that $E_{GS}=-N$, and the variance of the local energy cost scales as $1/\sqrt{d_{C}}$ as a result.
We emphasize that we do not guarantee that MSFO will always be able to find the ground state for any appropriately normalized MAX-3-XORSAT instance, given that it cannot overcome TFC. We presume that there are also additional hypergraph properties that ensure exponential scaling for MSFO (and our other filtered optimization methods), for both exact and approximate optimization; we just do not know what they are. If the transition to the ground state becomes first order and is missed, MSFO could likely suffer from the same field weakening and interference problems as TAQC, though this could be addressed by choosing an approximation target $A<1$, something that could be interesting to explore in future work.
VI.4 Discussion of our spectrally folded optimization results
@@
The reader may note that all of our spectral folding simulations targeted energies in the range $A=0.75$ to $A=0.85$, well above the theoretical prediction of $\sim 0.6$. We also tested $A=0.65$ at smaller scales (data not shown) for both formulations but found no region of the parameter space where it showed meaningful improvements over choosing $A=0.75$ or higher. When increasing $A$ to 0.85 we were not able to meet this target with constant probability, though for some cases we did see improvements in the returned achievable approximation ratio $q_{a}$ when compared to $A=0.75$. Aside from our MSFO simulations where $A=1$, we did not run tests with $A>0.85$ due to system size constraints, though comparing performance between $A=0.75$ and $A=0.85$ in table 1 suggests we have found the performance ceiling for these methods. Given the comparatively small system sizes available to classical simulation our ability to draw meaningful scaling distinctions with very small changes in $A$ is limited. Considering the results of all our numerical simulations, the minimum achievable approximation ratio for our PPSP problem classes is at least 0.7 (often 0.75) for the quadratic AQC variation, and at least 0.75 (often 0.8) for the 3-XORSAT TMA variation.
@@
As summarized in table 1, the contrast between the clear super-polynomial decay of higher approximation ratios with classical methods and TAQC (Figs. 6-9), and the constant probabilities returned by folded quantum optimization (Figs. 10-13) is stark. In particular, for the established methods, the polynomial time approximation threshold is set by the problem structure, specifically the relative energy of the high energy clustering phase, and as our PPSPs become more extremal this threshold steadily worsens as a fraction of the energy of the planted ground state. We found no evidence for an exponential separation in approximation power between local classical searches and TAQC. We assume, but did not test, that this would hold for linear-depth QAOA as well, though we do not rule out the possibility of meaningful speedups given fine tuning. At lower constraint densities, the clustering energy is not far above the ground state and spectrally folded quantum optimization provides no benefits for approximation, though it can still show scaling advantages for returning states close in Hamming distance to the ground state (see appendices). At higher constraint densities however, spectral folding is able to return states close to $G$ (and often find it directly, given quadratic runtimes and multiple stages), in both energy and Hamming distance, consistently and without degradation as $N_{C}/N$ increases. The performance of all spectral folding variations tested was broadly similar, with the 3-XORSAT TMA variation returning the highest approximation ratio but at the highest prefactor cost in gate count.
@@
This illustrates the fundamentally different structure of collective tunneling in this approach. Where TAQC and other direct methods attempt to find the ground state, asymptotically fail (given exponential scaling), and return approximate states passed after this missed opportunity, spectrally folded optimization deforms the cost function to search for states in an exponentially large hyperspherical shell, avoiding the interference and weakening field issues that are qualitatively responsible for TAQC’s lack of obvious quantum advantage. In the interest of fair comparisons the total runtime per-shot of all routines besides MSFO is $O\left(N\right)$ Hamiltonian calls. With this simple linear scaling we are able to ensure that any exponential scaling of the time to solution not visible from our figures or fits must involve very small exponents, though of course we cannot rule out such behavior on numerics alone. We thus do not see our numerical simulations as a precision scaling benchmark of folded quantum optimization, but rather a test of the basic veracity of our conservative analytical predictions in section IV.
Likewise, if our goal is to find ground states and not just good approximations thereof, it may well be the case that algorithms targeting $A<1$ are actually superior in the approximation-hard regime, even if MSFO with $A=1$ can find the ground state directly in sub-quadratic time! This is because those algorithms are capable of reliably finding states near the ground state, e.g. in its basin of attraction, and from there simple classical methods such as our quasi-greedy algorithm or simulated annealing can likely relax to the ground state quickly. If the polynomial runtime scaling to find approximate solutions is favorable (e.g. $O\left(N\right)$ vs. $O\left(N^{1.46}\right)$), then relaxing from an approximate solution may be more efficient. The choices explored in this paper likely only scratch the surface of what is possible in filtered optimization. We present all of these simulations–and new quantum algorithms–as constructive evidence for the fast approximability of MAX-3-XORSAT through a novel speedup mechanism. The results of our simulations, for a wide range of parameters, all suggest that our theory is sound.
VI.5 Extensions of these results
@@
Having thoroughly explored MAX-3-XORSAT in this work, it would be interesting to test spectral filtering methods on other hard CSPs or problems that can be straightforwardly formulated as such. We expect that the derivation of the achievable approximation ratio through the overlap of dressed states could generalize with some modifications to many other problem classes. We focused on MAX-3-XORSAT due to the simplicity of its structure, classical approximation hardness, and the fact that its exponential difficulty scaling is obvious at small $N$ for more standard classical and quantum approaches.
@@
One interesting potential exception to this argument is quantum Monte Carlo. Being a stoquastic problem, a folded quantum spin glass can in principle be efficiently simulated using QMC Isakov et al. (2016); Andriyash and Amin (2017); Jiang et al. (2017a, b); King et al. (2019), which for a uniform field has in some cases proven to be an effective quantum-inspired classical solver Heim et al. (2015); Albash and Lidar (2018). Incorporating spectral folding into these algorithms is possible, though the loss of locality makes evaluating each update much more expensive. As QMC is, fundamentally, an energy-based classical local update rule combined with many replicas, we do not expect it to overcome the entropic barrier in MAX-3-XORSAT the way that true quantum evolution can, though of course have not tested this here. In this vein, we note also the results of Hastings, and subsequently Gilyén and Vazirani, that a super-polynomial separation exists between stoquastic AQC (the class containing some of the algorithms in this work) and classical computation Hastings (2021); Gilyén et al. (2021); the stoquasticity of our algorithms thus does not rule out super-polynomial speedups. All that said, there may be some narrow cases where incorporating spectral filtering in a QMC calculation could prove useful as a classical solver; this would be an interesting avenue for future research.
If it turns out that many or all of these instances are efficiently approximable classically through QMC with a filtered spectrum (or any other classical method, for that matter), that would be a very significant discovery in its own right. Nothing in this work rules that out, and we think the issue deserves further inquiry.
@@ VII Conclusions and Outlook
Using the NP-hard MAX-3-XORSAT problem class, we explored the question of classical and quantum approximation hardness from a practical, mechanism-focused point of view. Guided by theoretical intuition, we proposed a class of instances called planted partial solution problems (PPSPs) which we showed are empirically hard for both exact and approximate optimization for classical searches and established quantum methods such as adiabatic quantum computing (AQC) and, we presume (but did not directly test), the quantum approximate optimization algorithm (QAOA). Through extrapolation and qualitative analysis of a rigorous calculation of the typical-case minimum gap at the first order paramagnet to spin glass transition bottlenecking MAX-3-XORSAT, we were able to identify two effects that significantly impede the ability of high-depth Trotterized AQC (TAQC) to approximate the hardest instances, namely weakening transverse field and destructive interference. We then proposed a novel algorithmic advance, called spectrally filtered quantum optimization, that provides multiple routes to circumvent these issues. The filtering transformation itself is conceptually simple; the analytical analysis of the full algorithm’s performance is perhaps less so. And rather intriguingly, it works by applying classical modifications to the classical cost function being optimized, which only provide benefits for quantum optimization.
@@
Using a resummed extensive-order perturbation theory, we were able to predict a constant fraction approximation guarantee for our difficult random hypergraph PPSPs, and consequently, an exponential quantum speedup in the classically hard regime. To further support our claims, we performed a series of numerical simulations of high-depth TAQC, classical optimization routines, and spectrally filtered optimization methods, out to the largest sizes that we could feasibly reach while still being able to gather good statistics. These numerical results support our claims and we did not discover any meaningful discrepancies or red flags in them; every major prediction here has been numerically checked to the extent we found reasonably possible.
@@
The implications of an effective fast approximation guarantee through spectrally filtered optimization–or any quantum algorithm, for that matter–are profound, and even restricting ourselves to MAX-3-XORSAT it is important to ask what types of hypergraph might cause it to fail. To be clear, “failure” in this case means that, for a class of hypergraphs, some property invalidates the analytical predictions we made and restricts polynomial-time approximation to values of $q_{m}$ near zero. Such a property, whether present in large random instances or specific to a tiny subclass thereof, would have to break both folded optimization through collective tunneling, and the softening of the first order phase transition in the multistage composition of folding and warping.
Focusing on the trial minimum annealing formulation, if a subsequent, more sophisticated analysis shows, for example, that due to some subtle effect missed in our resummed extensive order perturbation theory, the variations we propose are asymptotically limited to $q_{m}=0.4$ for random hypergraphs instead of our prediction of $\geq 0.6$, we would not consider that a general failure of the algorithm. Given the exponential density of target states, for folded quantum optimization to fail the per-state tunneling rate needs to be much worse than what our calculations and simulations return; even doubling the per-state tunneling decay exponent in Eq. (IV.3.2) yields $q_{m}\simeq 0.18$, a much worse approximation ratio than we predict and observe in simulations, but still a constant fraction better than random guessing. And this estimate is, again, for a formulation of our trial minimum annealing algorithm that did not use all the tools available to us, such as spectral warping, that would further improve performance. And finally, we emphasize again that neither of our two filtered optimization algorithms used all the techniques available to improve approximation performance–targeting excited states by folding in the case of multistage filtered optimization, and spectral warping in trial minimum annealing.
@@
We are not so hubristic as to claim it is impossible that there are hidden effects that substantially worsen scaling at large $N$ for random hypergraphs, which are not captured by our theories and are invisible in our numerics. But we see no evidence for it, have no idea what such an issue could be, and formulated our theory such that the simplifying approximations we made were more likely to underestimate the achievable approximation ratio $q_{a}$ than overestimate it. Excellent empirical performance in simulation, and the fact that two completely different calculations analyzing completely different filtered optimization algorithm formulations yield similar speedups, both bolster the strength of our claims. We see it as much more likely that one can structure PPSP hypergraphs in some non-random way as to violate the core assumptions of our calculations and become inapproximable. The conditions on such hypergraphs are, however, fairly strict; besides needing to more than double the exponential decay rate of per-state tunneling in folded optimization (as compared to random hypergraphs), whatever property is responsible for the slowdown must be resilient both to the addition of a random, uncorrelated problem with similar constraint density and ground state energy, and to spectral deformations such as nonlinear folding and spectral warping. Further, such graphs must be relatively dense, as sparse problems are easy to approximate by solving random sub-problems, and their construction rules must not inadvertently render them amenable to classical optimization.
@@
Identifying hypergraph properties that cause this entire class of methods to break down could lead to valuable new discoveries about macroscopic quantum tunneling physics and problem hardness, and further algorithm innovations in the steps needed to mitigate their effects. As we stated in the introduction, the underlying reasons for classical approximation hardness are generic and intuitive if often only applicable to extremal problem instances, but their quantum equivalents are not, and much more opaque. The correctness of our predictions here would imply that quantum approximation hardness may not be generic at all, and instead specific to as-yet undiscovered sets of problem properties. Let’s go exploring.